OpenStudy (leilaalemtz):

This is a physics question= it's about vectors (V1)= vector 1 If (V⃗ 1) and (V⃗ 2) are perpendicular, (V1)*(V2)=

OpenStudy (gabbyalicorn):

v^1v^2 =v^1*v^2 =(v)*(v*v) =v*v*v =v^3

OpenStudy (leilaalemtz):

Thank you but not what I'm looking :)

OpenStudy (anonymous):

opps, dot product makes the most sense

OpenStudy (michele_laino):

hint: scalar product between orthogonal vectors is equal to zero

OpenStudy (leilaalemtz):

it is dot product

OpenStudy (anonymous):

Do you recall a formula for dot product with an angle in it?

OpenStudy (leilaalemtz):

do you mean ABcos(angle)????

OpenStudy (anonymous):

\[a\cdot b = |a||b|cos\theta\]

OpenStudy (anonymous):

yep! what is cos of 90?

OpenStudy (anonymous):

perpendicular ---> right angles

OpenStudy (leilaalemtz):

I got it!! thank you!

OpenStudy (anonymous):

np